an interest rate of 14.90% per year compounded every 3 months is equivalent to a weekly compounded interest rate of

15.16 %

14.65%

14.88%

19.02%


1
Expert's answer
2022-03-17T06:47:14-0400

A=P(1+r/n)nt

Assume we have a principle of 50000, which is to be compounded over a ten year period


=50000(1+14.9%/4)40

=50000(1.03725)40

=50000(4.3185)


=215925


So we now are going to compute the rate of interest that is required to compound 50000 into 215925 over a ten year period.


215925=50000(1+r/52)520

215925/50000=(1+r/52)520

4.3185=(1+r/52)520

We now find the 520th root on both sides to get the below equation

1.002817=1+r/52

Solving it we find that the rate is 14.65%

R=14.65%


So the correct answer is 14.65%